# Convergence of the open WASEP stationary measure without Liggett’s condition

### v1

@@ Abstract.
We demonstrate that it is not necessary to assume Liggett’s condition in order to obtain convergence of the open ASEP stationary measures to the open KPZ stationary measure. This is equivalent to demonstrating that, under weak asymmetry scaling and appropriate scaling of time and space, the four-parameter Askey-Wilson process converges to a two-parameter continuous dual Hahn process. We conjecture that the convergence of the open ASEP height function process to solutions to the open KPZ equation will hold for a wider range of ASEP parameters than those permitted by Liggett’s condition.
@@ 1.1. Preface
The open ASEP is a relatively simple model of a system with nonzero net current in its equilibrium state, and, despite its simple definition, it exhibits complex behaviors associated to many-particle systems, such as phase transitions. Due to these properties, the open ASEP has been of great interest to researchers in probability and mathematical physics. Bertini and Giacomin [BG97] discovered that the ASEP on the full line converges to the Kardar-Parisi-Zhang (KPZ) equation on the full line under weak asymmetry scaling assumptions (see Section 1.3 for a longer discussion of this relationship). Corwin and Shen [CS18] then discovered that the Open ASEP on two additional domains (half-open and open ASEP) converge to the half-space and open KPZ equations, respectively, under triple point scaling of the boundary parameters and weak asymmetry scaling. Parekh [Par19] extended the results from Corwin and Shen which were restricted to a certain range of parameters of ASEP to hold for all parameters. As discovered by Bertini and Giacomin [BG97], the open ASEP is closely related to the Kardar-Parisi-Zhang (KPZ) equation (see Section 1.3 for a longer discussion of this relationship). Corwin and Knizel [CK24] used the connection between the open ASEP and the open KPZ equation to give the first construction of the stationary measure for the open KPZ equation through a scaling limit of open ASEP stationary measures. Work by Derrida, Evans, Hakim, and Pasquier [DEHP93] introduced an algebraic technique, called the matrix product ansatz (MPA) for obtaining the stationary measure of the open ASEP. Uchiyama, Sasamoto, and Wadati [USW04] used the MPA expression for the stationary measure of open ASEP to express the measure in terms of Askey-Wilson polynomials. More recently, Bryc and Wesolowski [BW17] demonstrated a formula for the multi-point correlation function of the stationary measure in terms of a process called the Askey-Wilson Process (which has one-point and transition densities given in terms of Askey-Wilson polynomials). Corwin and Knizel characterized the Laplace transform of the finite dimensional marginals of the stationary distribution of the open KPZ equation in terms of a stochastic process called the continuous dual Hahn process (see Section 2.3), which is the limit of the Askey-Wilson process under the appropriate scaling.
@@
The KPZ equation has a unique stationary measure, which was proved by Knizel and Matetski [KM22] for $u+v>0$ and $\min(u,v)>-1$, and by Parekh for all $u,v\in\mathbb{R}$. Corwin and Knizel characterized this stationary measure in terms of its Laplace transform for all $u+v\geq 0$. Barraquand and Le Doussal [BLD22] and Bryc, Kuznetsov, Wang, and Wesolowski [BKWW23] showed that the stationary measure of the open KPZ equation constructed by Corwin and Knizel, has a probabilistic description as a process $\{H_{u,v}(x)\}_{x\in[0,1]}=\{\mathbb{B}(\frac{x}{2})-\mathbb{X}_{u,v}(x)\}_{x\in[0,1]}$, where $\mathbb{B}(x)$ is standard Brownian motion and $\mathbb{X}_{u,v}(x)$ is an exponential re-weighting of Brownian motion by a Radon-Nikodym derivative
@@
Where $f\in C[0,1]$. For the purpose of this paper, we will use the description of the open ASEP in terms of its multipoint Laplace transform, see (3.2). Recently, work by Barraquand, Corwin, and Yang [BCY] constructed the KPZ stationary measure for all $u,v\in\mathbb{R}$ as a scaling limit from the log-gamma directed polymer model, modulo a result extending the existing understanding of convergence of the spacetime process of the log-gamma polymer to convergence on an interval $[0,L]$ (see Claim 4.2, [BCY]).
Corwin and Knizel’s construction of the open KPZ stationary measure assumed a relation, called Liggett’s condition, on the ASEP parameters. This relation was also used in earlier works by Corwin and Shen [CS18] and Parekh [Par19] that showed that the open ASEP height function process converges as a space-time process to solutions of the open KPZ. The contribution of this paper is to demonstrate that Liggett’s condition is not necessary to obtain convergence of the stationary measures of open ASEP to that of open KPZ. Based on this result, we conjecture that the convergence of open ASEP to open KPZ as a space-time process will also hold without this assumption.
This paper also demonstrates an interesting property of the Askey-Wilson process and its limiting process, the continuous dual Hahn process. The Askey-Wilson process is defined in general with parameters $(A,B,C,D,q)$ (see 2.1), but Liggett’s condition fixes the values of $B$ and $D$ as functions of $q$. The fact that the limit of the Askey-Wilson process when $B$ and $D$ are allowed to vary is still a continuous dual Hahn process which depends only on $A$ and $C$ is an interesting observation about the structure of the Askey-Wilson and continuous dual Hahn processes, which have been studied in works such as [BW10, BWW23, BW19, Bry22, WWY]. This observation was anticipated by Bryc, Wang, and Wesolowski [BWW23] (see, for example, the remarks before Assumption 1.1 in this paper).
@@ 1.2. The Open ASEP
The underlying lattice of the open ASEP has finite size $N$. We use $\alpha,\gamma$ to denote the rate of particle movement into and out of the leftmost lattice point from, respectively to, the left reservoir. Similarly, we use $\beta,\delta$ to denote the rate of particle movement out of, respectively into, the rightmost lattice point of the lattice to, respectively from, the right reservoir. Within the lattice, particles jump at times which are determined by independent, exponentially distributed random variables, with rate $1$ for movement to the right and rate $q\in(0,1)$ for movement to the left. The state space of the open ASEP is $\{0,1\}^{N}$. We refer to elements of the state space by $\tau=(\tau_{1}(t),...,\tau_{N}(t)$ where $\tau_{i}(t)=1$ if and only if there is a particle in position $i$ at time $t$. The open ASEP process $\tau(t)$ is a Markov process on the state space $\{0,1\}^{N}$ and the infinitesimal generator $\mathcal{L}(\cdot)$ which acts on functions of the open ASEP state space $f:\{0,1\}^{N}\to\mathbb{R}$ by
$(\mathcal{L}f)(\tau)$ $=\left(\alpha(1-\tau_{1})+\gamma\tau_{1}\right)\cdot\left(f(\tau^{1})-f(\tau)\right),$
$+\sum_{i=1}^{N-1}\left(\tau_{i}(1-\tau_{i+1})+q(1-\tau_{i})\tau_{i+1}\right)\cdot\left(f(\tau^{i,i+1})-f(\tau)\right),$
$+\left(\beta\tau_{N}+\delta(1-\tau_{N})\right)\cdot\left(f(\tau^{N})-f(\tau)\right).$
In this equation, the notation $\tau^{i}$ means that, in $\tau$, the value of $\tau_{i}$ is sent to $1-\tau_{i}$ and the notation $\tau^{i,i+1}$ means that $\tau_{i}$ and $\tau_{i+1}$ are swapped. We will denote the stationary measure of the open ASEP with parameters $\alpha,\beta,\gamma,\delta,q$ and lattice size $N$ as $\pi_{N}(\cdot)$, suppressing the dependence on the boundary parameters and the asymmetric jump rate. For the duration of this work, we will use $\langle f\rangle_{N}$ to denote the expectation of a function $f$ under $\pi_{N}(\cdot)$. The time evolution of the open ASEP induces a Markov process, called the height function process,
@@
where $h_{N;\alpha,\beta,\gamma,\delta}(t,0)$ is the net number of particles which have entered through the left boundary at time $t$. If $\tau$ is distributed according to $\pi_{N}(\cdot)$, we refer to the law of the process $x\mapsto h_{N;\alpha,\beta,\gamma,\delta}(t,x)-h_{N;\alpha,\beta,\gamma,\delta}(t,0)$ as the stationary measure for the open ASEP height function increment process.
@@
(1.1) $A=q^{v},$ $B=-q,$ $C=q^{u},$ $D=-q,$ $q=\exp{\left(\frac{-2}{\sqrt{N}}\right)},$
and a particular choice of scaling for the height function (see 1.4). This choice of scaling for $q$ is called weak asymmetry scaling. Corwin and Knizel showed that for any $u,v\in\mathbb{R}$, the $N$-indexed collection of open ASEP stationary measures, scaled according to the weak asymmetry scaling in $N$, was tight, and that sub-sequential limits were stationary measures of the open KPZ equation. Under the constraint that $u+v>0$, they also showed that all sub-sequential limits were equal, and gave a formula for the stationary measures in terms of their Laplace transforms. Later results by Knizel and Matetski [KM22] (for $u+v>0$ and $\min\{u,v\}>-1$) and and Parekh [Par] (for all $u,v\in\mathbb{R}$) demonstrated that the open KPZ stationary measure is unique.
The scaling of $A,B,C,D$ with respect to $q$ was chosen to preserve two properties. It preserves a $N^{-\frac{1}{2}}$ scaling around the triple point,
@@
(1.2) $\alpha+\gamma/q=1,$ $\beta+\delta/q=1.$
The contribution of this paper is to observe that, even without assuming Liggett’s condition, the scaling limit of the stationary measure of the open ASEP still converges to the stationary measure of the open KPZ equation. The new scaling assumptions are collected below.
Assumptions 1.1.
@@
(1.4) $\alpha=\frac{1-q}{(1-q^{r})(1+q^{u})},$ $\beta=\frac{1-q}{(1-q^{w})(1+q^{v})},$
@@
To demonstrate that the wider range of parameters captures a wider range of open ASEP stationary measures, we consider the case where $N=2$ and then separately compute the stationary measure in the case where $q=\frac{1}{2}$ and $u=v=1$, with $w=r=1$ and $w=2,r=3$.
Example 1.2.
Fixing $\rho_{\ell}=(1+q)^{-1},\rho_{r}=q(1+q)^{-1}$ (equivalently, $u=v=1$), we vary $w$ and $r$:
  1. (1)
    Setting $w=r=1$, we have $(\alpha,\beta,\gamma,\delta)=\left(\frac{2}{3},\frac{2}{3},\frac{1}{6},\frac{1}{6}\right)$. The stationary measure is given by
    ($\mathbb{P}(\{0,0\})=\frac{1}{4},$ ($\mathbb{P}(\{1,0\})=\frac{1}{4},$ ($\mathbb{P}(\{0,1\})=\frac{1}{4},$ ($\mathbb{P}(\{1,1\})=\frac{1}{4}.$
  2. (2)
    Setting $w=2,r=3$, we obtain $(\alpha,\beta,\gamma,\delta)=\left(\frac{8}{21},\frac{4}{9},\frac{1}{14},\frac{1}{18}\right)$ and stationary measure is given by
    ($\mathbb{P}(\{0,0\})$ $=\frac{35845251}{140695480}\approx 0.255$ ($\mathbb{P}(\{1,0\})=\frac{7259395}{28139096}\approx 0.258$
    ($\mathbb{P}(\{0,1\})$ $=\frac{6549543}{28139096}\approx 0.233$ ($\mathbb{P}(\{1,1\})=\frac{35805539}{140695480}\approx 0.254.$
This example illustrates that altering the parameters $w$ and $r$ effects the stationary measure of the open ASEP when $N$ is finite. The result of this paper demonstrates that these small differences disappear in the scaling limit, and the limiting stationary measure depends only on $u$ and $v$. It would be interesting to study how the rate of convergence of the open ASEP stationary measures varies in $w$ and $r$.
@@ 1.3. Statement and Discussion of the Theorem
where $\xi(t,x)$ is space-time white noise. Solutions of the open KPZ equation satisfy Neumann boundary conditions $(\partial_{x}H)(t,0)=u$ and $(\partial_{x}H)(t,1)=-v$. This definition is slightly informal. For background information on the open KPZ equation and a careful definition, we direct the reader to Corwin and Shen [CS18] or Parekh [Par19]. Both of those works proved convergence of the open ASEP height function process to Hopf-Cole solutions to the open KPZ equation for a range of boundary parameters under the scaling conditions in 1.4, below.
The unique open KPZ stationary measure is the law of the process $x\mapsto H_{0}(x)-H_{0}(0)$ on $(C[0,1],\mathcal{B}(C[0,1]))$ such that the law of the space-time process $(t,x)\mapsto H(t,x)-H(t,0)$, produced by evolving initial data $H(0,x)-H(0,0)$ (with $H_{0}(x)=H(0,x)$) according to the open KPZ equation, is invariant in time. Corwin and Knizel [CK24] gave an explicit formula (for $u+v\geq 0$) for the open KPZ stationary measure in terms of the Laplace transform of the continuous dual Hahn process (3.2). We will denote the stationary measure of the open KPZ equation with boundary parameters $u,v\in\mathbb{R}$ as the law of the process $x\mapsto H_{u,v}(x)-H_{u,v}(0)$ for $x\in[0,1]$. Using this notation, we can state the main result of this article.
@@
It is natural to consider which of these scaling assumptions is necessary to obtain the convergence of the open ASEP space-time process to that of the open KPZ. Convergence results for various ASEP models to KPZ variants by Corwin and Shen [CS18]; Parekh [Par19]; and Yang [Yan22] have followed the framework of a foundational result by Bertini and Giacomin [BG97], who approached the problem via the Hopf-Cole transform. Consequently, the argument in each of these works requires Liggett’s condition. While Gonçalves, Perkowski and Simon[GPS20] do not use the Hopf-Cole transform, they still assume Liggett’s condition. It is not known whether convergence as processes occurs under the full range of boundary parameters for the open ASEP. The effective boundary densities, $\rho_{\ell}$ and $\rho_{r}$, depend only on $A$ and $C$, which suggests we may vary $B$ and $D$ without disrupting the triple point scaling. Our result in 1.3 provides further evidence that it is possible to obtain convergence as space-time processes without Liggett’s condition, since it is possible at the level of stationary measures. We conjecture that convergence as space-time processes holds using the scaling assumptions in 1.1.
Assumptions 1.4.
In [CS18], the scaling assumptions for the convergence of the open ASEP height function $h_{N}(\cdot)$ to the Hopf-Cole solution to the open KPZ equation are
  1. (1)
    Weak asymmetry scaling: $q=\exp{\left(\frac{-2}{\sqrt{N}}\right)}$
@@
    $\alpha+\gamma/q=1$ $\beta+\delta/q=1$
@@
    $h^{(N)}_{u,v}(t,x)$ $:=N^{-\frac{1}{2}}h_{N;\alpha,\beta,\gamma,\delta}\left(\frac{1}{2}e^{N^{-\frac{1}{2}}}N^{2}t,Nx\right)+\left(\frac{1}{2}N^{-1}+\frac{1}{24}\right)t$
    $z^{(N)}_{u,v}(t,x)$ $:=\exp{\left(h_{u,v}^{(N)}(t,x)\right)}$
@@
    $\mathbb{E}\left[|z^{(N)}_{u,v}(x)|^{n}\right]^{1/n}\leq C(n)$ $\mathbb{E}\left[|z^{(N)}_{u,v}(x)-z^{(N)}_{u,v}(x^{\prime})|^{n}\right]^{1/n}\leq C(\theta,n)|x-x^{\prime}|^{\theta}$
@@
Conditions $(1)$ and $(3)$ are the same as the equivalently numbered conditions in 1.1 in the stationary case, but Liggett’s condition is an additional assumption.
@@
We denote the Hopf-Cole solution to the open KPZ equation with Neumann boundary parameters $u$ and $v$ by $H(t,x)$. We define $Z(t,x):=\exp{\left(H(t,x)\right)}$. We denote the diffusively scaled height function process with boundary rates $\alpha,\beta,\gamma,\delta$ given by (1.4) as $h_{u,v,w,r}^{(N)}(t,x)$, and $z_{u,v,w,r}^{(N)}(t,x):=\exp{\left(h_{u,v,w,r}^{(N)}(t,x)\right)}$. Under the scaling assumptions in 1.1, and using the height function process scaling defined in 1.4 (4) and requiring the same Hölder bounds described in 1.4 (5), then, if the initial data $z^{(N)}_{u,v,w,r}(0,x)$ converges to $Z(0,x)$, $z^{(N)}_{u,v,w,r}(t,x)\implies Z(t,x)$ as a space-time process.
1.4. Approach to proving 1.3
The proof of the tightness result in 1.3 is an application of an attractive coupling of ASEP proved in [CK24]. This result allows us to obtain Hölder bounds on $h^{(N)}_{u,v,w,r}(x)-h^{(N)}_{u,v,w,r}(0)$ which consequently allow us to conclude that the sequence is tight. For the details of this approach, see Section 5
In order to prove that the sequence of measures in 1.3 converges to the law of $H_{u,v}(\cdot)-H_{u,v}(0),$ we rely on the fact that the open ASEP stationary measure can be written in terms of a stochastic process called the Askey-Wilson process. Consequently, we can describe the Laplace transform of the height function process in terms of the Askey-Wilson process and analyze the asymptotics of that expression in terms of Askey-Wilson densities. We also rely on the fact that pointwise convergence of Laplace transforms on an open set implies convergence in distribution (see 2.6). The main technical contribution of this paper is to show that in a more general range of parameters, for any $w,r>0$ and $u+v>0$, the Askey-Wilson process in parameters $v,w,u,r$ (see Section 2.2.2) converges to the continuous dual Hahn process with parameters $u,v$. Consequently, when $u+v>0$ and $w,r>0$, the tight sequence of laws from 1.3 has a unique limit, which is the stationary measure of the open KPZ equation.

### v2

@@ Abstract.
We demonstrate that Liggett’s condition can be relaxed without disrupting the convergence of open ASEP stationary measures to the open KPZ stationary measure. This is equivalent to demonstrating that, under weak asymmetry scaling and appropriate scaling of time and space, the four-parameter Askey-Wilson process converges to a two-parameter continuous dual Hahn process. We conjecture that the convergence of the open ASEP height function process to solutions to the open KPZ equation will hold for a wider range of ASEP parameters than those permitted by Liggett’s condition.
@@ 1.1. Preface
The open ASEP is a relatively simple model of a system with nonzero net current in its equilibrium state, and, despite its simple definition, it exhibits complex behaviors associated to many-particle systems, such as phase transitions. Due to these properties, the open ASEP has been of great interest to researchers in probability and mathematical physics. Bertini and Giacomin [BG97] discovered that the ASEP on the full line converges to the Kardar-Parisi-Zhang (KPZ) equation on the full line under weak asymmetry scaling assumptions (see Section 1.3 for a longer discussion of this relationship). Corwin and Shen [CS18] then proved that the open ASEP on two additional domains (half-open and open ASEP) converge to the half-space and open KPZ equations, respectively, under triple point scaling of the boundary parameters and weak asymmetry scaling. Parekh [Par19] extended the results from [CS18] which were restricted to a certain range of parameters of ASEP to hold for all parameters. As discovered in [BG97], the open ASEP is closely related to the Kardar-Parisi-Zhang (KPZ) equation (see Section 1.3 for a longer discussion of this relationship). Corwin and Knizel [CK24] used the connection between the open ASEP and the open KPZ equation to give the first construction of the stationary measure for the open KPZ equation through a scaling limit of open ASEP stationary measures. Their work built on extensive literature about the open ASEP, developed over the preceding decades. Work by Derrida, Evans, Hakim, and Pasquier [DEHP93] introduced an algebraic technique, called the matrix product ansatz (MPA) for obtaining the stationary measure of the open ASEP. Uchiyama, Sasamoto, and Wadati [USW04] used the MPA expression for the stationary measure of open ASEP to express the measure in terms of Askey-Wilson polynomials. More recently, Bryc and Wesolowski [BW17] demonstrated a formula for the multi-point correlation function of the stationary measure in terms of a process called the Askey-Wilson Process (which has marginal and transitional laws determined by Askey-Wilson polynomials). [CK24] characterized the Laplace transform of the finite dimensional marginals of the stationary distribution of the open KPZ equation in terms of a stochastic process called the continuous dual Hahn process (see Section 2.3), which is the limit of the Askey-Wilson process under the scaling assumptions which take the open ASEP stationary measure to the stationary measure of the open KPZ equation (see 1.2).
@@
The open KPZ equation has a unique stationary measure, which was proved by Knizel and Matetski [KM22] for $u+v>0$ and $\min(u,v)>-1$, and by Parekh [Par22] for all $u,v\in\mathbb{R}$. [CK24] characterized this stationary measure in terms of its Laplace transform for all $u+v\geq 0$. Barraquand and Le Doussal [BLD22] and Bryc, Kuznetsov, Wang, and Wesolowski [BKWW23] showed that the stationary measure of the open KPZ equation constructed in [CK24] has a probabilistic description as a process $\{H_{u,v}(x)\}_{x\in[0,1]}=\{B(\frac{x}{2})-X_{u,v}(x)\}_{x\in[0,1]}$, where $B(x)$ is standard Brownian motion and the law of $X_{u,v}(x)$ is an exponential re-weighting of the law of a Brownian motion by a Radon-Nikodym derivative
@@
Where $f\in C[0,1]$. For the purpose of this paper, we will use the description of the open KPZ equation in terms of its multi-point Laplace transform, see (3.2). Recently, work by Barraquand, Corwin, and Yang [BCY23] constructed the KPZ stationary measure for all $u,v\in\mathbb{R}$ as a scaling limit from the log-gamma directed polymer model, modulo a result extending the existing understanding of convergence of the spacetime process of the log-gamma polymer to convergence on an interval $[0,L]$ (see Claim 4.2 [BCY23]).
The construction [CK24] of the open KPZ stationary measure assumed a relation, called Liggett’s condition, on the ASEP parameters. This relation was also used in earlier works [CS18, Par19] that showed that the open ASEP height function process converges as a space-time process to solutions of the open KPZ equation. The contribution of this paper is to demonstrate that Liggett’s condition is not necessary to obtain convergence of the stationary measures of open ASEP to that of open KPZ equation. Based on this result, we conjecture that the convergence of open ASEP to open KPZ equation as a space-time process will also hold without this assumption.
This paper also demonstrates an interesting property of the Askey-Wilson process and its limiting process, the continuous dual Hahn process. The Askey-Wilson process is defined in general with parameters $(A,B,C,D,q)$ (2.1), but Liggett’s condition fixes the values of $B$ and $D$ as functions of $q$. It is an interesting observation about the structure of the Askey-Wilson and continuous dual Hahn processes, which have been studied in works such as [BW10, BWW23, BW19, Bry22, WWY23], to note that the parameters $B$ and $D$ have no effect on the limit process. This observation was anticipated by Bryc, Wang, and Wesolowski [BWW23] (see, for example, the remarks before Assumption 1.1 in this paper).
@@ 1.2. The Open ASEP
The underlying lattice of the open ASEP has finite size $N$. We will now define the Markov process of particle movement on this lattice in terms of transition rates. We use $\alpha,\gamma$ to denote the rate of particle movement into and out of the leftmost lattice point from, respectively to, the left reservoir. Similarly, we use $\beta,\delta$ to denote the rate of particle movement out of, respectively into, the rightmost lattice point of the lattice to, respectively from, the right reservoir. Within the lattice, particles jump at times which are determined by independent, exponentially distributed random variables, with rate $1$ for movement to the right and rate $q\in(0,1)$ for movement to the left. The state space of the open ASEP is $\{0,1\}^{N}$. We refer to elements of the state space by $\tau=(\tau_{1},...,\tau_{N})$ and use the notation $\tau_{i}(t)$ to denote the occupancy of lattice point $i$ at time $t$, with $\tau_{i}(t)=1$ if and only if there is a particle in position $i$ at time $t$. The open ASEP process $\tau(t)$ is a Markov process on the state space $\{0,1\}^{N}$ and the infinitesimal generator $\mathcal{L}(\cdot)$ which acts on functions of the open ASEP state space $f:\{0,1\}^{N}\to\mathbb{R}$ by
$(\mathcal{L}f)(\tau)$ $=\left(\alpha(1-\tau_{1})+\gamma\tau_{1}\right)\cdot\left(f(\tau^{(1)})-f(\tau)\right)$
$+\sum_{i=1}^{N-1}\left(\tau_{i}(1-\tau_{i+1})+q(1-\tau_{i})\tau_{i+1}\right)\cdot\left(f(\tau^{(i,i+1)})-f(\tau)\right)$
$+\left(\beta\tau_{N}+\delta(1-\tau_{N})\right)\cdot\left(f(\tau^{(N)})-f(\tau)\right).$
In this equation, the notation $\tau^{(i)}$ means that, in $\tau$, the value of $\tau_{i}$ is sent to $1-\tau_{i}$ and the notation $\tau^{(i,i+1)}$ means that $\tau_{i}$ and $\tau_{i+1}$ are swapped. We will denote the stationary measure of the open ASEP with parameters $\alpha,\beta,\gamma,\delta,q$ and lattice size $N$ as $\pi_{N}(\cdot)$, suppressing the dependence on the boundary parameters and the asymmetric jump rate. For the duration of this work, we will use $\langle f\rangle_{N}$ to denote the expectation of a function $f$ under $\pi_{N}(\cdot)$. The time evolution of the open ASEP induces a Markov process, called the height function process,
@@
where $h_{N;\alpha,\beta,\gamma,\delta}(t,0)$ is the net number of particles which have entered through the left boundary at time $t$. If $\tau$ is distributed according to $\pi_{N}(\cdot)$, we refer to the law of the process $\{h_{N;\alpha,\beta,\gamma,\delta}(t,x)-h_{N;\alpha,\beta,\gamma,\delta}(t,0)\}_{x\in[0,1]}$ as the stationary measure for the increment process of the open ASEP height function. In this case the law is independent of $t$, and we will also denote the process by $\{h_{N;\alpha,\beta,\gamma,\delta}(x)-h_{N;\alpha,\beta,\gamma,\delta}(0)\}_{x\in[0,1]}$.
@@
$A=q^{v},$ $B=-q,$ $C=q^{u},$ $D=-q,$ $q=\exp{\left(\frac{-2}{\sqrt{N}}\right)},$ (1.1)
and a particular choice of scaling for the height function (1.2). This choice of scaling for $q$ is called weak asymmetry scaling. [CK24] showed that for any $u,v\in\mathbb{R}$, the $N$-indexed collection of open ASEP stationary measures, scaled according to the weak asymmetry scaling in $N$, was tight, and that sub-sequential limits were stationary measures of the open KPZ equation. Under the constraint that $u+v>0$, they also showed that all sub-sequential limits were equal, and gave a formula for the stationary measures in terms of their Laplace transforms. Later results [KM22, Par22] demonstrated that the open KPZ stationary measure is unique.
The scaling of $A,B,C,D$ with respect to $q$ was chosen to preserve two properties. It preserves a $N^{-\frac{1}{2}}$ scaling around the triple point (discussed in [Cor22]),
@@
$\alpha+\gamma/q=1,$ $\beta+\delta/q=1.$ (1.2)
The contribution of this paper is to observe that, without assuming Liggett’s condition, the scaling limit of the stationary measure of the open ASEP still converges to the stationary measure of the open KPZ equation. The new scaling assumptions are collected below.
Assumption 1.1.
@@
$\alpha=\frac{1-q}{(1-q^{r})(1+q^{u})},$ $\beta=\frac{1-q}{(1-q^{w})(1+q^{v})},$ (1.4)
@@
In the equation above, imposing Liggett’s condition would set $w=r=1$ and result in a factor of $(1-q)$ canceling from the numerator and denominator in each expression.
@@ 1.3. Statement and Discussion of the Theorem
where $\xi(t,x)$ is space-time white noise. Solutions of the open KPZ equation satisfy Neumann boundary conditions $(\partial_{x}H)(t,0)=u$ and $(\partial_{x}H)(t,1)=-v$. This definition is slightly informal. For background information on the open KPZ equation and a careful definition, we direct the reader to [CS18, Par19]. Both of those works proved convergence of the open ASEP height function process to Hopf-Cole solutions to the open KPZ equation for a range of boundary parameters under the scaling conditions in 1.2, below.
The unique open KPZ stationary measure is the law of the process $\{H_{0}(x)-H_{0}(0)\}_{x\in[0,1]}$ on the space $(D[0,1],\mathcal{B}(D[0,1]))$ such that the law of the space-time process $\{H(t,x)-H(t,0)\}_{t\in\mathbb{R}_{\geq 0},x\in[0,1]}$, produced by evolving initial data $H(0,x)-H(0,0)$ (with $H_{0}(x)=H(0,x)$) according to the open KPZ equation, is invariant in time. [CK24] gave an explicit formula (for $u+v\geq 0$) for the open KPZ stationary measure in terms of the Laplace transform of the continuous dual Hahn process (3.2). We will denote the stationary measure of the open KPZ equation with boundary parameters $u,v\in\mathbb{R}$ as the law of the process $\{H_{u,v}(x)-H_{u,v}(0)\}_{x\in[0,1]}$. Using this notation, we can state the main result of this article.
@@
It is natural to consider which of these scaling assumptions is necessary to obtain the convergence of the open ASEP space-time process to that of the open KPZ equation. Convergence results for various ASEP models to KPZ variants [CS18, Par19, Yan22] have followed the framework of the foundational paper [BG97], which approached the problem via the Hopf-Cole transform. Consequently, the argument in each of these works requires Liggett’s condition. While Gonçalves, Perkowski, and Simon [GPS20] do not use the Hopf-Cole transform, they still assume Liggett’s condition. It is not known whether convergence as processes occurs for any choice of $u,v,w,r$ satisfying $u+v>0$ and $w,r>0$. The effective boundary densities, $\rho_{\ell}$ and $\rho_{r}$, depend only on $A$ and $C$, which suggests we may vary $B$ and $D$ without disrupting the triple point scaling. Our result in 1.1 provides further evidence that it is possible to obtain convergence as space-time processes without Liggett’s condition, since it is possible at the level of stationary measures. We conjecture that convergence as space-time processes holds using the scaling assumptions in 1.1.
Assumption 1.2.
The scaling assumptions for the convergence of the open ASEP height function $h_{N}(\cdot)$ to the Hopf-Cole solution to the open KPZ equation are given by condition $(1)$ (weak asymmetry scaling) from 1.1, with the added conditions
@@
    $\alpha+\gamma/q=1,$ $\beta+\delta/q=1.$
@@
    $h^{(N)}_{u,v}(t,x)$ $:=N^{-\frac{1}{2}}h_{N;\alpha,\beta,\gamma,\delta}\left(\frac{1}{2}e^{N^{-\frac{1}{2}}}N^{2}t,Nx\right)+\left(\frac{1}{2}N^{-1}+\frac{1}{24}\right)t,$
    $z^{(N)}_{u,v}(t,x)$ $:=\exp{\left(h_{u,v}^{(N)}(t,x)\right)}.$
@@
    $\mathbb{E}\left[|z^{(N)}_{u,v}(x)|^{n}\right]^{1/n}\leq C(n),$ $\mathbb{E}\left[|z^{(N)}_{u,v}(x)-z^{(N)}_{u,v}(x^{\prime})|^{n}\right]^{1/n}\leq C(\theta,n)|x-x^{\prime}|^{\theta},$
@@
We denote the Hopf-Cole solution to the open KPZ equation with Neumann boundary parameters $u$ and $v$ by $H_{u,v}(t,x)$. We define $Z_{u,v}(t,x):=\exp{\left(H_{u,v}(t,x)\right)}$. We denote the diffusively scaled height function process with boundary rates $\alpha,\beta,\gamma,\delta$ given by (1.4) as $h_{u,v,w,r}^{(N)}(t,x)$, and $z_{u,v,w,r}^{(N)}(t,x):=\exp{\left(h_{u,v,w,r}^{(N)}(t,x)\right)}$. Under weak asymmetry scaling assumptions (1.1), and using the height function process scaling defined in 1.2 (4) and requiring the same Hölder bounds described in 1.2 (5), then, if the initial data $z^{(N)}_{u,v,w,r}(0,x)$ converges to $Z_{u,v}(0,x)$, $z^{(N)}_{u,v,w,r}(t,x)\implies Z_{u,v}(t,x)$ as a space-time process.
1.4. Approach to proving 1.1
The proof of the tightness result in 1.1 is an application of an attractive coupling of ASEP proved in [CK24]. This result allows us to obtain Hölder bounds on $h^{(N)}_{u,v,w,r}(x)-h^{(N)}_{u,v,w,r}(0)$ which consequently allow us to conclude that the sequence is tight. For the details of this approach, see Section 5.
In order to prove that the sequence of measures in 1.1 converges to the law of $H_{u,v}(\cdot)-H_{u,v}(0),$ we rely on the fact that the open ASEP stationary measure can be written in terms of a stochastic process called the Askey-Wilson process. Consequently, we can describe the Laplace transform of the height function process in terms of the Askey-Wilson process and analyze the asymptotics of that expression in terms of Askey-Wilson densities. We also rely on the fact that pointwise convergence of Laplace transforms on an open set implies convergence in distribution (see 2.4). The main technical contribution of this paper is to show that in a more general range of parameters, for any $w,r>0$ and $u+v>0$, the Askey-Wilson process tangent process in parameters $v,w,u,r$ (see Section 2.2.2) converges under 1.1 to the continuous dual Hahn process with parameters $u,v$. Consequently, when $u+v>0$ and $w,r>0$, the tight sequence of laws from 1.1 has a unique limit, which is the stationary measure of the open KPZ equation.